What is compressibility?

What is rigidity?
What is fluidity?
What is density?
What is kinetic energy?
1
Expert's answer
2011-06-24T07:55:39-0400
1. In thermodynamics and fluid mechanics, compressibility is a measure of the relative volume change of a fluid or solid as a response to a pressure (or mean stress) change. Let V be a volume and p be a pressure. Then compressibility, denoted by β is defined as follows:
β = -1/V * V'p

2. Rigidity is a measure of the momentum of the particle.
It is defined as
R = B ρ = p/q,
where
* B is the magnetic field,
* ρ is the gyroradius of the particle due to this field,
* p is the particle momentum,
* and q is its charge.
It is frequently referred to as simply B ρ.

3. Fluidity is the inverse to viscosity.
The latter is a measure of the resistance of a fluid which is being deformed by either shear stress or tensile stress.
Suppose he fluid moves under the forse F.
Then the applied force, F, is proportional to the area, A, and velocity gradient, u, in the fluid and inversely proportional to the distance, y, between the plates:
F = μ A u / y.
The coeficient μ is called viscosity.
Then fluidity is
φ = 1 /μ.
4. The density of a material is defined as its mass per unit volume:
ρ = m / V
5. The kinetic energy of an object is the energy which it possesses due to its motion. If an object of mass m moves with velocity v, then its kinetic energy is
m * v2 / 2